A woman on a bridge 85.4 m high sees a raft floating at a constant speed on the river below. She drops a stone from rest in an attempt to hit the raft. The stone is released when the raft has 5.58 m more to travel before passing under the bridge. The stone hits the water 3.35 m in front of the raft. Find the speed of the raft.
Answer #1

Given hieght of bridge (h ) = 85.4 m

For free fall motion we have formula :

h = 1/2gt^2

g = 9.8 m/s^2

Putting in formula,

t = √(2×85.4/9.8) = 4.17 s

Distance traveled by raft in this time interval is :

d = initial position of raft – final position. = 5.58 – 3.35

d = 2.23 m

As raft travelling with constant speed we have formula ,

v = d/t = 2.23/4.17

speed of raft (v) = 0.53 m/s

​​​​​​Thank you please rate the solution.